he pressure and temperature at the beginning of compression of a cold air-standard Diesel cycle are 100 kPa and 300 K, respectively. At the end of the heat addition, the pressure is 7.2 MPa and the temperature is 2250 K. Assume constant specific heats evaluated at 300 K.


Determine:

(a) the compression ratio.

(b) the cutoff ratio.

(c) the percent thermal efficiency of the cycle.

(d) the mean effective pressure, in kPa.
Answer #1

Solution - Ti = 300k, P = 100 kpa, = 7.2M Pa = 7200 kPa Pe = Pz = 7200 kPa. T3 = 2250k Also K = 1.4 + T. = ( ) ET (200)*= 300Thermal efferiency is : n = lalnet Qin Man = 4 (T3-T₂) - Co (TG-T,) C G (T3-T2) = 1.005( 2250-1018.09) -0.748 (910-300) 1.005